# Fuchs’ problem for linear groups

### v1

@@ 1. Introduction
    If $|F|$ is odd and congruent to ±$\pm 1$ modulo 8, then $\mathbf{SL}_{2}(F)$ is not realizable.
@@
    If $|F|$ is odd and congruent to ±$\pm 3$ modulo 8, then $\mathbf{SL}_{2}(F)$ is not the group of units in a finite ring.
@@
    The order of $\mathbf{Hol}(\mathbf{Z}_{n})$ is a divisor of $48$
@@ 2.1. Finite rings
Our group $G$, of order 24, is also known as the binary tetrahedral group. The quaternion group $\mathbf{Q}_{8}=\{\pm 1,\pm i,\pm j,\pm k\}$ is a proper normal subgroup of $G$, and $G$ is generated by $\mathbf{Q}_{8}$ and an element $c\in G\setminus\mathbf{Q}_{8}$ of order 3. (Note that in the group ring $\mathbf{Z}_{t}[G]$, the element $-1\in\mathbf{Q}_{8}$ of order 2 is not the same as the additive inverse of the multiplicative identity $1\in R$.)
First, we consider the case $t=2$. We used SageMath/GAP to verify that $A=1+cj+(-c)$ and $B=j+c+ci$ are units in $\mathbf{Z}_{2}[G]$ of order dividing 8. Hence, there are elements $x,y\in\mathbf{Q}_{8}$ such that $A+x$ and $B+y$ are elements of $I$. However, again using SageMath/GAP, we checked that in all cases the natural map $G\longrightarrow\mathbf{Z}_{2}[G]/(A+x,B+y)$ has a nontrivial kernel, contradicting the fact that the natural map $G\longrightarrow\mathbf{Z}_{2}[G]/I$ is an embedding.
For the case $t=4$, we proceed similarly. Suppose $R=\mathbf{Z}_{4}[G]/I$ realizes $G$ and the natural map $G\longrightarrow\mathbf{Z}_{4}[G]/I$ is an embedding. Since $i^{2}$ is the unique element of order 2 in $G=R^{\times}$, we must have $01+i^{2}=$ 0 in $R$. Thus, $1+i^{2}\in I$. Since $(i+j+k)^{2}=1,$ we have that $i+j+k\pm 1\in I.$ For any $r\in G$, $(1+2r)^{2}=1$, so either $02r=$ 0 or $22r=$ 2 mod $I$. But since $r$ is a unit and $2\neq 0$ in $R,$ we must have $22r=$ 2 mod $I$ for all $r\in G$. Now, let $x=1+c+i.$ Then,
@@ 2.2. Rings of characteristic zero
    If $q\equiv 1\,(4),$ then $\mathbf{UC}(\mathbf{F}_{q})$ is cyclic of order $q-1$.
@@
from which it follows that $z=-y$ and $x=w$, so $A\in\mathbf{UC}(\mathbf{F}_{q}).$ Hence, $\mathbf{UC}(\mathbf{F}_{q})$ is a maximal abelian subgroup of $\mathbf{SL}_{2}(\mathbf{F}_{q})$. ∎
@@
Assume to the contrary that there is a ring $R$ of characteristic zero whose group of units is isomorphic to $\mathbf{SL}_{2}(\mathbf{F}_{q}).$ Then, the maximal abelian subgroup $\mathbf{UC}(\mathbf{F}_{q})$ is also realizable in characteristic zero. If $q\equiv 1\,(8)$, then $q\equiv 1\,(4)$ and $\mathbf{UC}(\mathbf{F}_{q})=\mathbf{C}_{q-1}$ with $18\mid q-$ 1 . If $q\equiv-1\,(8),$ then $q\equiv 3\,(4)$ and $\mathbf{UC}(\mathbf{F}_{q})=\mathbf{C}_{q+1}$ with $18\mid q+$ 1. In both cases, we may conclude that a cyclic group whose order is divisible by 8 is realizable in characteristic zero; however, this is impossible by [PS70]. ∎
@@ 3. Affine linear groups
    If $R$ is a ring with group of units $\mathbf{Hol}(\mathbf{Z}_{n})$, the the characteristic of $R$ is $0,2,3,4,$ or $6$; if $n$ is odd, then $R$ has characteristic $2$
@@
These relations force $b=1$ and $02a=$ 0. The center is therefore trivial if $n$ is odd and cyclic of order 2, generated by $(n/2,1)$, if $n$ is even.
@@ 3.1. Finite rings
For the even primary case, assume that $G=\mathbf{Hol}(\mathbf{Z}_{2^{k}})$ is isomorphic to a dihedral group. As above, there is some cyclic $H\leq G$ with index $2$. If $\mathbf{Z}_{2^{k}}\leq H$, then we get that $\varphi(2^{k})=2$ as above, in which case $42^{k}=$ 4. If instead we have that $\mathbf{Z}_{2^{k}}\not\leq H$, then there must be some generator $g$ of $\mathbf{Z}_{2^{k}}$ such that $g\in\mathbf{Z}_{2^{k}}\setminus H$. Because $g$ is a generator of $\mathbf{Z}_{2^{k}}$, it necessarily has order $2^{k}$; however, since $G$ is dihedral and $g$ lies outside of $H$, the order of $g$ must be $2$, and so $22^{k}=$ 2.
Finally, we prove that (A) and (E) are equivalent. Observe that $|\mathbf{Hol}(\mathbf{Z}_{n})|=n\,\varphi(n)$. Simple calculations show that $n\,\varphi(n)$ is a divisor of $48$ for all $n\mid 12$. Because $\varphi(n)$ is even for all $n\geq 3$, if $n\,\varphi(n)\mid 48$ it follows that $n\mid 24$. Note that every divisor of $24$ except for $8$ and $24$ are divisors of $12$, thus the equivalence of (A) and (E) now follows from the fact that $328\varphi(8)=$ 32 and $19224\varphi(24)=$ 192 do not divide $48$. ∎

### v2

@@ 1. Introduction
    If $|F|$ is odd and congruent to $\pm 1$ modulo 8, then $\mathbf{SL}_{2}(F)$ is not realizable.
@@
    If $|F|$ is odd and congruent to $\pm 3$ modulo 8, then $\mathbf{SL}_{2}(F)$ is not the group of units in a finite ring.
@@
    The order of $\mathbf{Hol}(\mathbf{Z}_{n})$ is a divisor of $48.$
@@ 2.1. Finite rings
Our group $G$, of order 24, is also known as the binary tetrahedral group. The quaternion group $\mathbf{Q}_{8}=\{\pm 1,\pm i,\pm j,\pm k\}$ is a proper normal subgroup of $G$, and $G$ is generated by $\mathbf{Q}_{8}$ and an element $c\in G\setminus\mathbf{Q}_{8}$ of order 3. (Note that in the group ring $\mathbf{Z}_{t}[G]$, the element $1\in\mathbf{Q}_{8}$ of order 2 is not the same as the additive inverse of the multiplicative identity $1\in R$.)
First, we consider the case $t=2$. We used SageMath/GAP to verify that A=1+cj+(c)A=1+cj+(-c) and $B=j+c+ci$ are units in $\mathbf{Z}_{2}[G]$ of order dividing 8. Hence, there are elements $x,y\in\mathbf{Q}_{8}$ such that $A+x$ and $B+y$ are elements of $I$. However, again using SageMath/GAP, we checked that in all cases the natural map $G\longrightarrow\mathbf{Z}_{2}[G]/(A+x,B+y)$ has a nontrivial kernel, contradicting the fact that the natural map $G\longrightarrow\mathbf{Z}_{2}[G]/I$ is an embedding.
For the case $t=4$, we proceed similarly. Suppose $R=\mathbf{Z}_{4}[G]/I$ realizes $G$ and the natural map $G\longrightarrow\mathbf{Z}_{4}[G]/I$ is an embedding. Since $i^{2}$ is the unique element of order 2 in $G=R^{\times}$, we must have $1+i^{2}=0$ in $R$. Thus, $1+i^{2}\in I$. Since $(i+j+k)^{2}=1,$ we have that $i+j+k\pm 1\in I.$ For any $r\in G$, $(1+2r)^{2}=1$, so either $2r=0$ or $2r=2$ mod $I$. But since $r$ is a unit and $2\neq 0$ in $R,$ we must have $2r=2$ mod $I$ for all $r\in G$. Now, let $x=1+c+i.$ Then,
@@ 2.2. Rings of characteristic zero
    If $q\equiv 1\,(4),$ then $\mathbf{UC}(\mathbf{F}_{q})$ is cyclic of order q1q-1.
@@
from which it follows that z=yz=-y and $x=w$, so $A\in\mathbf{UC}(\mathbf{F}_{q}).$ Hence, $\mathbf{UC}(\mathbf{F}_{q})$ is a maximal abelian subgroup of $\mathbf{SL}_{2}(\mathbf{F}_{q})$. ∎
@@
Assume to the contrary that there is a ring $R$ of characteristic zero whose group of units is isomorphic to $\mathbf{SL}_{2}(\mathbf{F}_{q}).$ Then, the maximal abelian subgroup $\mathbf{UC}(\mathbf{F}_{q})$ is also realizable in characteristic zero. If $q\equiv 1\,(8)$, then $q\equiv 1\,(4)$ and $\mathbf{UC}(\mathbf{F}_{q})=\mathbf{C}_{q-1}$ with $8\mid q-1.$ If $q\equiv-1\,(8),$ then $q\equiv 3\,(4)$ and $\mathbf{UC}(\mathbf{F}_{q})=\mathbf{C}_{q+1}$ with $8\mid q+1$. In both cases, we may conclude that a cyclic group whose order is divisible by 8 is realizable in characteristic zero; however, this is impossible by [PS70]. ∎
@@ 3. Affine linear groups
    If $R$ is a ring with group of units $\mathbf{Hol}(\mathbf{Z}_{n})$, the the characteristic of $R$ is $0,2,3,4,$ or $6$; if $n$ is odd, then $R$ has characteristic $2.$
@@
These relations force $b=1$ and $2a=0$. The center is therefore trivial if $n$ is odd and cyclic of order 2, generated by $(n/2,1)$, if $n$ is even.
@@ 3.1. Finite rings
For the even primary case, assume that $G=\mathbf{Hol}(\mathbf{Z}_{2^{k}})$ is isomorphic to a dihedral group. As above, there is some cyclic $H\leq G$ with index $2$. If $\mathbf{Z}_{2^{k}}\leq H$, then we get that $\varphi(2^{k})=2$ as above, in which case $2^{k}=4$. If instead we have that $\mathbf{Z}_{2^{k}}\not\leq H$, then there must be some generator $g$ of $\mathbf{Z}_{2^{k}}$ such that $g\in\mathbf{Z}_{2^{k}}\setminus H$. Because $g$ is a generator of $\mathbf{Z}_{2^{k}}$, it necessarily has order $2^{k}$; however, since $G$ is dihedral and $g$ lies outside of $H$, the order of $g$ must be $2$, and so $2^{k}=2$.
Finally, we prove that (A) and (E) are equivalent. Observe that $|\mathbf{Hol}(\mathbf{Z}_{n})|=n\,\varphi(n)$. Simple calculations show that $n\,\varphi(n)$ is a divisor of $48$ for all $n\mid 12$. Because $\varphi(n)$ is even for all $n\geq 3$, if $n\,\varphi(n)\mid 48$ it follows that $n\mid 24$. Note that every divisor of $24$ except for $8$ and $24$ are divisors of $12$, thus the equivalence of (A) and (E) now follows from the fact that $8\varphi(8)=32$ and $24\varphi(24)=192$ do not divide $48$. ∎